OpenStudy (chrisga):

Given the system of equations presented here: 4x + y = 4 2x + 7y = 28 Which of the following actions creates an equivalent system such that, when combined with the other equation, one of the variables is eliminated?

OpenStudy (chrisga):

Multiply the second equation by −1 to get −2x − 7y = −28 Multiply the second equation by −4 to get −8x − 28y = −112 Multiply the first equation by −7 to get −28x − 7y = −28 Multiply the first equation by −2 to get −8x − 2y = −8

OpenStudy (chrisga):

someone please help me to answer this will fan and medal